OpenStudy (anonymous):

i have 3 questions since i already finish but i don't want to fail it. Please make sure your right and I'm right as well. 1.) solve for x 4(x-b)+x A.)b=4/3x B.)b=3/4x C.)x=4/3b D.)x=3/4b <--- 2.) solve for y, then find the value of y when given x=-2 6x=7-4y A.)-12<---- B.)7/4 C.)19/4 D.)24 3.) solve for H A=1/2bh A.)2a/b=h b.)2a=b/h<---- c.)h=1/2ab h=2ab All of the answers i pick have a arrow pointing at them and the ones with a / between them mean fraction.

OpenStudy (rhystic):

2 is wrong should be 19/4 3 is wrong should be h = 2a/b

OpenStudy (rhystic):

1 is correct by the way

OpenStudy (rhystic):

work for 2 : 6*-2 = -12 -12 = 7 - 4y -12-7 = -4y -19 = -4y 19/4 = y work for 3: a = bh/2 2a = bh 2a/b = h

OpenStudy (anonymous):

thank u very much